# Roman Garipov $^{*\,\dagger}$

### v2

@@ Abstract
We introduce AutoJudge [^1] Our code is available at github.com/garipovroma/autojudge., a method that accelerates large language model (LLM) inference with task-specific lossy speculative decoding. Instead of matching the original model output distribution token-by-token, we identify which of the generated tokens affect the downstream quality of the response, relaxing the distribution match guarantee so that the “unimportant” tokens can be generated faster. Our approach relies on a semi‑greedy search algorithm to test which of the mismatches between target and draft models should be corrected to preserve quality and which ones may be skipped. We then train a lightweight classifier based on existing LLM embeddings to predict, at inference time, which mismatching tokens can be safely accepted without compromising the final answer quality. We evaluate the effectiveness of AutoJudge with multiple draft/target model pairs on mathematical reasoning and programming benchmarks, achieving significant speedups at the cost of a minor accuracy reduction. Notably, on GSM8k with the Llama 3.1 70B target model, our approach achieves up to ${\approx}2{\times}$ speedup over speculative decoding at the cost of ${\leq}1\%$ drop in accuracy. When applied to the LiveCodeBench benchmark, AutoJudge automatically detects programming-specific important tokens, accepting ${\geq}25$ tokens per speculation cycle at $2\%$ drop in Pass@1. Our approach requires no human annotation and is easy to integrate with modern LLM inference frameworks.
@@ 1 Introduction
A popular way to speed up LLM inference is through speculative decoding (Leviathan et al., 2023; Chen et al., 2023), which uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. This method, along with its successors (Miao et al., 2023; Cai et al., 2024; Li et al., 2024b), can speed up LLM inference while guaranteeing that the generated outputs match those of the original model (for greedy inference) or follow the same distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
@@
Our experiments with Llama 3.x models demonstrate that the proposed approach can indeed identify important tokens and save time on speculation, accepting on average over 40 tokens per target model forward pass, (approx. $2{\times}$ that of speculative decoding) at the cost ${\leq}1{\%}$ drop in accuracy on GSM8K (Cobbe et al., 2021) and even more with minor accuracy drawdown. We obtain similar results with the Qwen2.5(Yang et al., 2024) family models, observing comparable number of accepted tokens and accuracy trade-offs. When applied to programming tasks on LiveCodeBench (Jain et al., 2024), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms. The main contributions of our work can be summarized as follows:
  • We formulate AutoJudge, an algorithm for detecting which of the tokens generated in speculative decoding affect the downstream quality of a given task. Our algorithm requires no human annotation and can be applied to most popular LLM tasks.
  • We verify the efficacy of AutoJudge on mathematical reasoning and programming benchmarks for several speculative decoding setups (model pairs) for both mathematical reasoning and programming. Our evaluations demonstrate favorable tradeoffs between accuracy and inference speedup, generating 20-45 tokens per speculative decoding cycle at the cost of slight accuracy drawdown.
  • We integrate our inference procedure with the vLLM library (Kwon et al., 2023) and report inference speed on A100 GPUs for both 8B and 70B target models, with up to $2{\times}$ speedup over speculative decoding at ${\leq}1\%$ quality drawdown, and on H100 GPUs with a 405B target model.
@@ 2 Background
Our work builds on top of speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate K>1K{>}1 possible future tokens, then runs the main “target” model in parallel to verify∗*∗**For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches the sampling probabilities (Leviathan et al., 2023). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model. Subsequent works improve on this idea by generating draft trees instead of single sequences (Miao et al., 2023; Liu et al., 2023; Chen et al., 2024; Svirschevski et al., 2024), training specialized “heads” to draft next tokens based on the model’s hidden states (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024b, a), and more (Fu et al., 2023; Spector and Re, 2023; Sun et al., 2023; He et al., 2023).
@@ 3.1 Mining Important Tokens
Without the loss of generality, we also assume that there is a problem-specific way to extract the final answer from the model’s response, $a=\textsc{extractAnswer}(y)$. In mathematical reasoning tasks such as GSM8K Cobbe et al. (2021), the final answer is literally whatever the model puts after "the final answer (is)". In programming tasks, the “answer” would be the output from the testing system given the generated code — either a report about passing and failing tests or a testing error (e.g., an Out Of Memory or Syntax Error). Finally, we say that two answers are equivalent $a_{ref}{\equiv}a_{alt}$ if they are the same from the downstream task perspective. Note that this does not require them to be exactly equal as strings: in math problems, $\texttt{1.5}{\equiv}\texttt{3/2}$, whereas in programming tasks, two programs can be equivalent despite having different variable names. If the task at hand does not have a formalized evaluation procedure, e.g., general conversation agents, we can define $\textsc{extractAnswer}(y)=y$ and detect if two answers are equivalent using an LLM or human judges.
@@ 3.2 Classifier Training
Once we gather a dataset of task-specific important tokens with Alg. 1, we can train a classifier that would detect such tokens for use during inference. This classifier can, in principle, be any type of model, from a simple linear model or decision tree to a fine-tuned transformer layer. However, in our work, we default to training lightweight linear models with existing LLM hidden states as features, since those would introduce the least overhead during inference. There are several important design choices that can affect the effectiveness of such a classifier: we address each one separately.
@@ 3.3 Inference
Generality. In our initial experiments, we focus on traditional speculative decoding (Leviathan et al., 2023; Chen et al., 2023) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (Miao et al., 2023; Svirschevski et al., 2024; Chen et al., 2024) and single-model multi-head algorithms (Cai et al., 2024; Li et al., 2024b, a). This also means that our approach can be integrated into existing inference frameworks such as vLLM (Kwon et al., 2023), TensorRT-LLM (NVIDIA, 2023) or TGI (Hugging Face, 2023).
@@
Comparison with Judge Decoding. As we discussed earlier, our approach can be seen as an extension of Judge Decoding that enables automatic dataset mining. As such, the dataset generation algorithm from Section 3.1 can be used in conjunction with the Judge Decoding training and inference protocol, which appears to be similar to ours up to possible minor details. It would be interesting to compare the two dataset collection strategies directly. Unfortunately, the annotated dataset and the source code of Judge Decoding are not available, which makes such comparison difficult.
@@ 4 Experiments
Figure 3: Accuracy and the average number of accepted tokens on GSM8K for (left) 8-shot Llama-3.2 1B draft / Llama-3.1 8B target and (middle) 0-shot Llama 3.1 8B draft / Llama 3.1 70B target and (right) 8-shot Qwen2.5-0.5B draft / Qwen2.5-7.5B target. All models are Instruct versions.
We organize our evaluations as follows: in Section 4.1 we evaluate AutoJudge on the GSM8K (Cobbe et al., 2021) mathematical reasoning benchmark. Next, in Section 4.2, we evaluate our approach on programming tasks from LiveCodeBench (Jain et al., 2024). We use Llama 3.x model family across all setups, focusing on two model pairs: 1) Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target∗*∗**The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models, and the 3.1 version does not have the smaller 1B models. and 2) Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target. Additionally, we report results with Qwen2.5-0.5B (draft) / Qwen2.5-7.5B (target) on the GSM8K benchmark to showcase the method’s transferability across model families. Finally, Section 4.3 contains GPU inference speed benchmarks with our vLLM implementation where we consider Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target and Llama-3.1-8B-Instruct draft / Llama-3.1-405B-Instruct target setups. The 405B setup is omitted in Section 4.1 and Section 4.2, since those sections rely on transformers-based inference, which is significantly slower than vLLM. We run AutoJudge on top of the standard speculative decoding algorithm (Leviathan et al., 2023). Our main experiments run in bfloat16 precision (see Appendix C for details). We provide results for additional model pairs in Appendix D, extended inference benchmarks under various setups in Appendix F, evaluations with different window sizes in Appendix G, and inference with offloading experiments in Appendix H.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school mathematical problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters (Gao et al., 2021) for zero-shot and 8-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Alg. 1) if the extracted final answers (numbers) are equal. For reference, we provide an example of important token assignments found by our algorithm in Figure 2.
@@
The results in Figure 3 demonstrate that AutoJudge decoding can achieve substantial speedups. Notably, our algorithm can accept over 40 tokens per target model forward pass in 8-shot evaluations for the 8B draft / 70B target model pair with a ${\leq}1\%$ change in accuracy. Varying the classifier threshold allows us to achieve even greater speedups at the cost of several percentage points drop in accuracy. The heuristic-based top-$K$ baseline also achieves some speedups, but at the cost of a significantly higher accuracy drawdown. We report setups in Appendix D.1.
@@ 4.2 Programming with LiveCodeBench
Next, we test if the AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model family on the LiveCodeBench (Jain et al., 2024) programming benchmark. Here, we use the code_generation_lite∗*∗**huggingface.co/datasets/livecodebench/code_generation_lite dataset with the version tag release_v5. The dataset contains 880 programming tasks; we evaluate on all three subsets: easy, medium and hard. Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s built-in test suite. We otherwise follow the standard evaluation protocol for this benchmark and report Pass@1 in zero-shot setting.
@@ 4.3 Inference Benchmarks with vLLM
In this section, we report the GPU inference speed of AutoJudge classifiers integrated into vLLM speculative decoding. For these benchmarks, we use the same GSM8K evaluation setup as in Section 4.1 and evaluate 8B/70B and 8B/405B model pairs. We integrate AutoJudge classifier into vLLM’s built-in lossless speculative decoding based on Leviathan et al. (2023). We compare AutoJudge decoding against lossless speculative decoding with draft window size set to 32: for each configuration, we report the absolute speed (tokens per second), as well as the relative speedup with respect to each baseline. We run 8B/70B benchmarks on 4 A100-SXM GPUs with NVSwitch and 8B/405B on 8 H100-SXM5 80GB GPUs with NVSwitch. We provide configuration details in Appendix E.
Our results, reported in Table 1 demonstrate that the advantage of our algorithm in terms of accepted tokens (Sections 4.1 & 4.2) translates proportionally to real-world tokens per second. Namely, if we allow ${\leq}2\%$ change in accuracy, our vLLM inference with AutoJudge classifier can achieve $108.5$ tokens per second for 8B/70B model pair, which translates into up to $2.21\times$ speedup compared to regular speculative decoding. For 8B/405B model pair we achieve $56.1$ tokens per second. In Appendix F we report additional benchmark configurations with similar performance gains. Also, we provide evaluations with individually tuned window sizes in Appendix  G as well as evidence of vLLM inference inconsistency in Appendix  I.
Limitations.
Our approach assumes that the downstream task has a way to determine whether or not two solutions are equivalent. While this is true for many tasks, it is less obvious for tasks such as story writing, image editing or general conversation. While it is possible to run AutoJudge on these tasks with an LLM judge to determine if the two answers are equivalent, we did not test this in our experiments. Additionally, running Algorithm 1 consumes compute resources, mostly LLM inference, in the form of local inference or API calls. We discuss this in more detail in Appendix J.
Table 1: Inference benchmarks (with vLLM) of AutoJudge classifiers on the GSM8K dataset for (left) Llama 3.1 8B draft / 3.1 70B target and (right) Llama 3.1 8B draft / 3.1 405B target models. Bolded columns highlight configurations that achieve strong speedups with minimal (1-2%) accuracy degradation.
Llama 3.1 8B draft / 3.1 70B target
@@
Speed, tokens/s 63.8 94.6 108.5 108.0
Speculative Decoding:   49.2 tokens/s
Speedup(ours) 1.30 1.92 2.21 2.20
Llama 3.1 8B draft / 3.1 405B target
Threshold 0.01 0.04 0.07 0.14
Accuracy, % 96.7 94.9 93.4 90.7
Speed, tokens/s 47.7 56.1 59.2 61.5
Speculative Decoding:   44.5 tokens/s
Speedup(ours) 1.07 1.26 1.33 1.38

### v3

@@ Abstract
[…]
@@ 1 Introduction
A popular way to speed up LLM inference is through speculative decoding (Leviathan et al., 2023; Chen et al., 2023), which uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. This method, along with its successors (Miao et al., 2023; Cai et al., 2024; Li et al., 2024d), can speed up LLM inference while guaranteeing that the generated outputs match those of the original model (for greedy inference) or follow the same distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
@@
Our experiments with Llama 3.x models demonstrate that the proposed approach can indeed identify important tokens and save time on speculation, accepting on average over 40 tokens per target model forward pass (approx. $2{\times}$ that of speculative decoding), at the cost of a ${\leq}1{\%}$ drop in accuracy on GSM8K (Cobbe et al., 2021) and even more with a minor accuracy drawdown. We obtain similar results with the Qwen2.5(Yang et al., 2024) family of models, observing comparable number of accepted tokens and accuracy trade-offs. When applied to programming tasks on LiveCodeBench (Jain et al., 2024), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms. The main contributions of our work can be summarized as follows:
  • We formulate AutoJudge, an algorithm for detecting which of the tokens generated in speculative decoding affect the downstream accuracy for a given task. Our algorithm requires no human annotation and can be applied to most popular LLM tasks.
  • We verify the efficacy of AutoJudge on mathematical reasoning and programming benchmarks for several speculative decoding setups (model pairs). Our evaluations demonstrate favorable tradeoffs between the accuracy and the inference speedup, generating 20–45 tokens per speculative decoding cycle at the cost of a slight accuracy drawdown.
  • We integrate AutoJudge with the vLLM framework (Kwon et al., 2023) and report the inference speed on A100 GPUs for both 8B and 70B target models, with up to $2{\times}$ speedup over speculative decoding at a ${\leq}1\%$ quality decrease, and on H100 GPUs with a 405B target model.
@@ 2 Background
Our work builds on top of speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate K>1K{>}1 possible future tokens, then runs the main “target” model in parallel to verify∗*∗**For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches the sampling probabilities (Leviathan et al., 2023). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model. Subsequent works improve on this idea by generating draft trees instead of single sequences (Miao et al., 2023; Liu et al., 2023; Chen et al., 2024; Svirschevski et al., 2024), training specialized “heads” to draft next tokens based on the model’s hidden states (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024d, c), and more (Fu et al., 2023; Spector and Re, 2023; Sun et al., 2023; He et al., 2023).
@@ 3.1 Mining Important Tokens
Without the loss of generality, we also assume that there is a problem-specific way to extract the final answer from the model’s response, $a=\textsc{extractAnswer}(y)$. In mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021), the final answer is literally whatever the model puts after "the final answer (is)". In programming tasks, the “answer” would be the output from the testing system given the generated code — either a report about passing and failing tests or a testing error (e.g., an Out Of Memory or Syntax Error). Finally, we say that two answers are equivalent $a_{ref}{\equiv}a_{alt}$ if they are the same from the downstream task perspective. Note that this does not require them to be exactly equal as strings: in math problems, $\texttt{1.5}{\equiv}\texttt{3/2}$, whereas in programming tasks, two programs can be equivalent despite having different variable names. If the task at hand does not have a formalized evaluation procedure, e.g., general conversation agents, we can define $\textsc{extractAnswer}(y)=y$ and detect if two answers are equivalent using an LLM or human judges.
@@ 3.2 Classifier Training
Once we gather a dataset of task-specific important tokens with Algorithm˜1, we can train a classifier that would detect such tokens for use during inference. This classifier can, in principle, be any type of model, from a simple linear model or decision tree to a fine-tuned transformer layer. However, in our work, we default to training lightweight linear models with existing LLM hidden states as features, since those would introduce the least overhead during inference. There are several important design choices that can affect the effectiveness of such a classifier: we address each one separately.
@@ 3.3 Inference
Generality. In our initial experiments, we focus on traditional speculative decoding (Leviathan et al., 2023; Chen et al., 2023) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (Miao et al., 2023; Svirschevski et al., 2024; Chen et al., 2024) and single-model multi-head algorithms (Cai et al., 2024; Li et al., 2024d, c). This also means that our approach can be integrated into existing inference frameworks such as vLLM (Kwon et al., 2023), TensorRT-LLM (NVIDIA, 2023) or TGI (Hugging Face, 2023).
@@
Comparison with Judge Decoding. As we discussed earlier, our approach can be seen as an extension of Judge Decoding that enables automatic dataset mining. As such, the dataset generation algorithm from Section 3.1 can be used in conjunction with the Judge Decoding training and inference protocol, which appears to be similar to ours up to possible minor details. In Appendix N, we additionally compare against manual human annotation similar to Judge Decoding. Unfortunately, the original source code and data from Judge Decoding are not available, making it difficult to compare directly.
@@ 4 Experiments
Figure 3: Accuracy and the number of accepted tokens on GSM8K for (left) 8-shot Llama-3.2 1B draft / Llama-3.1 8B target and (right) 0-shot Llama 3.1 8B draft / Llama 3.1 70B target (all Instruct)
We organize our evaluations as follows: in Section 4.1 we evaluate AutoJudge on the GSM8K (Cobbe et al., 2021) mathematical reasoning benchmark. Next, in Section 4.2, we evaluate our approach on programming tasks from LiveCodeBench (Jain et al., 2024). Finally, Section 4.3 contains GPU inference speed benchmarks with our vLLM implementation. We focus on two pairs of Llama 3.x models: 1) Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target∗*∗**The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models, and the 3.1 version does not have the smaller 1B models. and 2) Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target. We also report results with Qwen2.5 models on the GSM8K benchmark to showcase the method’s transferability across model families. Our main experiments run in bfloat16 precision (see Appendix C). We run AutoJudge on top of the standard speculative decoding algorithm (Leviathan et al., 2023) in main experiments and explore EAGLE-2 in Appendix J.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school mathematical problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters (Gao et al., 2021) for zero-shot and 8-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Algorithm˜1) if the extracted final answers (numbers) are equal. For reference, we provide an example of important token assignments found by our algorithm in Figure 2.
@@
Figure 4: Accuracy and the average number of accepted tokens on GSM8K 8-shot for (left) Qwen2.5-0.5B draft / Qwen2.5-7B target and (right) Qwen2.5-0.5B draft / Qwen2.5-32B target (all Instruct).
@@
The results in Figure 3 demonstrate that AutoJudge decoding can achieve substantial speedups. Notably, our algorithm can accept over 40 tokens per target model forward pass in 8-shot evaluations for the 8B draft / 70B target model pair with a ${\leq}1\%$ change in accuracy. Varying the classifier threshold allows us to achieve even greater speedups at the cost of several percentage points drop in accuracy. The heuristic-based top-$K$ baseline also achieves some speedups, but at the cost of a significantly higher accuracy drawdown. We report additional setups in Appendix D.1.
@@ 4.2 Programming with LiveCodeBench
Next, we test if the AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model family on the LiveCodeBench (Jain et al., 2024) programming benchmark. Here, we use the code_generation_lite∗*∗**huggingface.co/datasets/livecodebench/code_generation_lite dataset with the version tag release_v5. The dataset contains 880 programming tasks; we evaluate on all three subsets: easy, medium and hard. Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s built-in test suite. We follow the standard evaluation protocol for this benchmark and report Pass@1 in zero-shot setting.
@@ 4.3 Inference Benchmarks with vLLM
Table 1: Inference speed benchmarks on GSM8K 0-shot with vLLM for (left) 1B draft / 8B target models with tuned window size (baseline = 8, AutoJudge = 10) and (right) for 8B draft / 70B target models (all Instruct) with tuned window size (baseline = 8, AutoJudge = 32 ).
Llama 3.2 1B draft / 3.1 8B target (0-shot)
Threshold 0.06 0.12 0.15 0.16
Accuracy, % 83.1 80.2 79.8 77.4
Speed, tokens/s 149.2 169.2 171.2 173.9
Speculative Decoding:   147.7 tokens/s
Speedup(ours) 1.01 1.14 1.15 1.17
Llama 3.1 8B draft / 3.1 70B target (0-shot)
@@
Speed, tokens/s 72.3 80.6 107.4 109.6
Speculative Decoding:   72.3 tokens/s
Speedup(ours) 1.0 1.11 1.49 1.52
In this section, we report the GPU inference speed of speculative decoding with AutoJudge classifiers. To benchmark real-world inference speed, we integrated AutoJudge into vLLM speculative decoding. We use the same GSM8K evaluation setup as in Section 4.1 and evaluate 3 model pairs from Llama 3.x family: 1B/8B, 8B/70B and 8B/405B draft/target respectively (all Instruct).
We compare AutoJudge decoding against lossless vLLM speculative decoding (Leviathan et al., 2023). To ensure fair comparison, we tune draft window size for AutoJudge and baseline independently (see Appendix G). In each configuration, we report the absolute speed in tokens per second (batch size 1), as well as the relative speedup over speculative decoding. We run 1B/8B model pair on a single A100-SXM4-80GB GPU; 8B/70B on 4 A100-SXM4-80G GPUs in tensor-parallel mode. Finally, the 8B/405B runs on 8 H100-SXM5-80GB GPUs with the 405B model loaded in FP8 precision. Additionally, we consider a setup where 8B/70B model pair runs on a single A100-SXM4-80GB GPU with RAM offloading (see Appendix H). We provide additional configuration details in Appendix E.
Table 2: Inference speed with vLLM for (left) Llama 3.1 8B draft / 405B target models on GSM8K 0-shot with tuned window size (baseline=14, AutoJudge=20). (right) Llama 3.1 8B draft / 70B target (all Instruct) on GSM8K 8-shot with offloading, tuned window size (baseline=10, AutoJudge=48).
Llama 3.1 8B draft / 3.1 405B target (0-shot)
Threshold 0.01 0.05 0.09 0.14
Accuracy, % 96.1 93.4 92.5 91.5
Speed, tokens/s 50.6 58.0 58.5 60.1
Speculative Decoding:   50.0 tokens/s
Speedup(ours) 1.01 1.16 1.17 1.20
Llama 3.1 8B draft / 3.1 70B target (offload)
Threshold 0.03 0.05 0.11 0.28
Accuracy, % 95.4 94.8 93.4 90.4
Speed, tokens/s 1.4 1.6 1.9 2.4
Speculative Decoding:   1.19 tokens/s
Speedup(ours) 1.20 1.31 1.59 1.98
Our results in Tables 1 and 2 demonstrate consistent improvements across all model pairs, with particularly high speedups for 8B/70B with and without offloading. This confirms that our earlier results in terms of accepted tokens (Sections 4.1 & 4.2) translate to real-world tokens per second. Namely, if we allow $\approx 2\%$ change in accuracy, our vLLM inference with AutoJudge classifier can achieve $107.4$ tokens per second for 8B/70B model pair, which translates into about $1.5\times$ speedup compared to regular speculative decoding. This improvement is even more noticeable for hybrid setup with offloaded target model (in Table 2), where AutoJudge achieves up to $2\times$ speedup.
Additional evaluations.
To better explore the wide variety of tasks, methods and hardware setups, we report multiple series of additional experiments in supplementary materials. We evaluate AutoJudge decoding with EAGLE-2 in Appendix J. Next, we generalize to open-ended problems (question answering and creative writing) with LLM-as-a-judge evaluation in Appendix L. We also evaluate how AutoJudge classifiers transfer to adjacent problems in Appendix M. Finally, we benchmark with equal window size in Appendix D and control for vLLM nondeterminism in Appendix  I.
Limitations.
Our approach assumes that the downstream task has a way to determine whether or not two solutions are equivalent. While this is true for many tasks, we found that AutoJudge sometimes struggles in open-ended tasks with no clear criteria for correct answers. It would be interesting to explore alternative algorithm designs more suitable for open-ended problems. Additionally, running Algorithm 1 consumes compute resources, mostly LLM inference, in the form of local inference or API calls. We discuss this in more detail in Appendices P and Q.